OpenStudy (anonymous):

If y varies inversely as x, and y = 3 when x = 6, find the value of x when y = 18 possible answers- 36, 6, 9, 1

sam (.sam.):

y=k/x

sam (.sam.):

3=k/6 k=18 --------------------- y=18/x 18=18/x x=?

OpenStudy (anonymous):

haha i got it thank you :)